OpenStudy (anonymous):

Can someone help me solve this. ----> 2 - 7i

OpenStudy (uri):

how do u want to solve it?

OpenStudy (uri):

into polar form?

OpenStudy (anonymous):

yes,sorry, for some reason you comments wasnt showing up.

OpenStudy (uri):

ok so |z| = (2)^2+(7)^2

OpenStudy (uri):

you find the radius first.

OpenStudy (uri):

then you calculuate the arguement

OpenStudy (uri):

theta = y/x => tan^(-1)(7/2)

OpenStudy (uri):

so polar form is \[ z= r(cos\theta + isin\theta\]

OpenStudy (uri):

Do you understand..?

OpenStudy (anonymous):

A little , Ill try to solve